### v1

# Nuclear matter properties from chiral-scale effective theory including a dilatonic scalar meson

Abstract
Chiral effective theory has become a powerful tool for studying the low-energy properties of QCD. In this work, we apply an extended chiral effective theory—chiral-scale effective theory—including a dilatonic scalar meson to study nuclear matter and find that the properties around saturation density can be well reproduced. Compared to the traditionally used Walecka-type models in nuclear matter studies, our approach improves the behavior of symmetry energy and the incompressibility coefficient in describing empirical data without introducing additional freedoms. Moreover, the predicted neutron star structures fall within the constraints of GW170817, PSR J0740+6620, and PSR J0030+0451, while the maximum neutron star mass can reach about ${}3M_{\odot}$ with a pure hadronic phase. Additionally, we find that symmetry patterns of the effective theory significantly impact neutron star structures. We believe that introducing this type of theory into nuclear matter studies can lead to a deeper understanding of QCD, nuclear matter, and compact astrophysical objects.
I Introduction
Study on nuclear matter (NM) has long been crucial for understanding both nuclear force and neutron star (NS) structures (see, e.g., Refs. Serot and Walecka (1986); Brown and Rho (2002); Fukushima and Hatsuda (2011); Lattimer and Prakash (2016); Baym et al. (2018); Ma and Rho (2020a); Brandes and Weise (2024) and references therein). The properties of NM are highly sensitive to details of nucleon interactions. A popular and widely used approach to describe these interactions is the one-boson-exchange (OBE) model and its variants, such as the Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989). These models typically involve $\pi$, $\sigma$, $\omega$ and $\rho$ meson exchanges, covering the effective range of nucleon forces from $0.5\ \mathrm{fm}$ to $2\ \mathrm{fm}$. Utilizing these models, NM properties can be calculated using the relativistic mean field (RMF) approach Walecka (1974); Wiringa et al. (1988), which is the most practical and economical framework to introduce density effects.
It is recognized that Walecka-type models lack considerations of QCD symmetry patterns, the valid region of effective operators, and theoretical errors. Chiral effective field theory ($\rm{\chi EFT}$), thanks to the pioneering works by Weinberg Weinberg (1979, 1991), offers a powerful framework for studying nuclear forces at long ranges, anchored on QCD symmetry. To develop a realistic model of nuclear forces, it is accepted that vector mesons—$\rho$ and $\omega$— and isoscalar-scalar meson $\sigma$ are indispensable. The former can be regarded as gauge fields of hidden local symmetry (HLS) in $\rm{\chi EFT}$  Bando et al. (1985, 1988); Harada and Yamawaki (2003). However, the inclusion of the $\sigma$ meson as an independent degree of freedom in $\rm{\chi EFT}$ is not straightforward, since the fourth scalar component of the chiral four-vector is integrated out when transitioning from the linear sigma model to the nonlinear sigma model—the leading term of $\rm{\chi EFT}$. Considering that the lowest-lying scalar meson $\sigma$ has a mass roughly equivalent to that of the kaon, which is a Nambu-Goldstone (NG) boson of chiral symmetry breaking in three-flavor $\rm{\chi EFT}$ ($\chi\rm EFT_{3}$), by supposing that QCD has a nonperturbative infrared fixed point, Crewther and Tunstall Crewther and Tunstall (2015); Catà et al. (2019) proposed that the lowest-lying scalar meson can be considered as the NG boson (dilaton) of the spontaneous breaking of scale symmetry. Consequently, $\rm{\chi EFT}$ is extended to include the scalar meson, resulting in the scale-chiral effective theory ($\chi\rm EFT_{\sigma}$).
Based on the terminologies of $\chi\rm EFT_{\sigma}$ and HLS, the $\chi\rm EFT$ with baryons is constructed in Refs. Lee et al. (2015); Li et al. (2017) at the leading chiral order to discuss nucleon interactions. It is denoted as bsHLS, with ‘b’ for baryon and ‘s’ for scale. In bsHLS, the potential from meson exchanges in OBE models can be reproduced by expanding bsHLS to the first order of linear field couplings but with additional symmetry considerations. By using the low-momentum potential $V_{\rm lowk}$ renormalization group approach Bogner et al. (2003); Holt et al. (2007), taking the “leading order scale symmetry (LOSS)” approximation where the trace anomaly effect enters only through the dilaton potential, which breaks scale symmetry explicitly and spontaneously, we found that the chiral-scale EFT with few parameters can successfully describe not only NM at the saturation density but also the compact-star matter at $n=(5-7)n_{0}$ Ma and Rho (2020a). A novel phenomena, which was not realized, is that in nuclear matter at densities $n=(2-4)n_{0}$ the sound velocity of NM saturates the conformal limit $v_{s}^{2}=1/3$, but the trace of the energy-momentum tensor does not vanish, that is, the NM exhibits a pseudo-conformal structure Ma and Rho (2019a, b); Ma et al. (2019).
In this work, we study NM properties using bsHLS with the RMF method. The corrections to LOSS in the baryonic part, which have been found crucial for understanding the quenched $g_{A}$ value in the super-allowed Gamow-Teller transitions of heavy nuclei Lubos et al. (2019); Ma and Rho (2020b). After identifying the NM properties around $n_{0}$, it was found that bsHLS using the RMF method can yield reasonable results. The data can be reproduced with the choice of $\beta^{\prime}=\frac{\partial\beta}{\partial\alpha}\sim 1$, consistent with the results of Refs. Ma and Rho (2018); Shao and Ma (2022), where a dilaton limit fixed point is assumed in the medium. The combination of $f_{\chi}m_{\sigma}$ is constrained to be $\simeq 2.3\times 10^{5}$ MeV2 which is consistent with what estimated before in the skyrmion crystal approach such that physically interesting results can be obtained Lee et al. (2003); Park et al. (2008). By comparing the symmetry energy $E_{\rm sym}$ and incompressibility coefficient $K(n)$ obtained from our bsHLS and Walecka-type models, we find that the bsHLS can make the symmetry energy stiff at subsaturation density but soft at intermediate densities to meet the constraints of GW170817 Abbott et al. (2018) and the neutron skin thickness of ${}^{208}\rm Pb$ Reed et al. (2021) simultaneously, without introducing additional freedoms such as the $\delta$ meson in Refs. Li et al. (2022); Ma and Ma (2024). In addition, we find that the incompressibility coefficient of bsHLS surges at intermediate densities, while the Walecka-type models exhibit gentler behaviors, resulting in a better description of NS structures using bsHLS. The maximum mass of NS can nearly reach $3M_{\odot}$ with a pure hadron phase, meeting the constraints of GW170817 Abbott et al. (2018, 2019) and PSR J0740+6620 Miller et al. (2021), whereas Walecka-type models can only reach $2M_{\odot}$ within these constraints, as analyzed in Ref. Guo et al. (2024). We find that this is due to the kink behavior of $\sigma$ expectations at intermediate densities, induced by the nonlinear realization of scale symmetry. Moreover, the value of $\beta^{\prime}$, behavior of $\langle\chi\rangle^{*}$ and Brown-Rho scaling (B-R scaling) at different densities can significantly affect NS structures, indicating a relationship between QCD symmetry patterns and macroscopic phenomena.
This article is organized as follows: In Sec. II, we introduce the theoretical framework of bsHLS and the equation of state (EoS) of NM under the RMF approximation. In Sec. III, we provide a phenomenological analysis by pinning the experimental data of nuclear matter and neutron star observations. A comparison with the Walecka-type models is also made. Our summary and discussion are presented in Sec. IV.
II bsHLS in nuclear medium
In this section, we establish the theoretical framework of bsHLS and the EoS of NM from the RMF approach. Here, we only interest in NM made of nucleons, we reduce the bsHLS Lagrangian to the two-flavor case, $u$ and $d$. The leading order bsHLS Lagrangian $\mathcal{L}$ can be decomposed into the baryonic part $\mathcal{L}_{B}$ and the mesonic part $\mathcal{L}_{M}$:
$\mathcal{L}=\mathcal{L}_{M}+\mathcal{L}_{B}\ .$ (1)
The mesonic part $\mathcal{L}_{M}$, consisting of $\sigma$, $\rho$, $\omega$ and $\pi$ mesons, is constructed as follows Li et al. (2017),
$\mathcal{L}_{M}$ $=$ $\ f^{2}_{\pi}\Phi^{2}\left[h_{1}+\left(1-h_{1}\right)\Phi^{\beta^{\prime}}\right]{\rm Tr}\left(\hat{\alpha}^{\mu}_{\perp}\hat{\alpha}_{\mu\perp}\right)$
${}+\frac{m^{2}_{\rho}}{g^{2}_{\rho}}\Phi^{2}\left[h_{2}+(1-h_{2})\Phi^{\beta^{\prime}}\right]{\rm Tr}\left(\hat{\alpha}^{\mu}_{\parallel}\hat{\alpha}_{\mu\parallel}\right)$
${}+\frac{1}{2}\left(\frac{m^{2}_{\omega}}{g^{2}_{\omega}}-\frac{m^{2}_{\rho}}{g^{2}_{\rho}}\right)\Phi^{2}\left[h_{3}+\left(1-h_{3}\right)\Phi^{\beta^{\prime}}\right]{\rm Tr}\left(\hat{\alpha}^{\mu}_{\parallel}\right){\rm Tr}\left(\hat{\alpha}_{\mu\parallel}\right)$
${}+\frac{1}{2}\left[h_{4}+\left(1-h_{4}\right)\Phi^{\beta^{\prime}}\right]\partial_{\mu}\chi\partial^{\mu}\chi+\frac{f^{2}_{\pi}}{4}\Phi^{3-\gamma_{m}}{\rm Tr}\left(\mathcal{M}U^{\dagger}+U\mathcal{M}^{\dagger}\right)+h_{5}\Phi^{4}+h_{6}\Phi^{4+\beta^{\prime}}$
${}-\frac{1}{2g^{2}_{\rho}}\left[h_{7}+\left(1-h_{7}\right)\Phi^{\beta^{\prime}}\right]{\rm Tr}\left(V_{\mu\nu}V^{\mu\nu}\right)-\frac{1}{2g^{2}_{0}}\left[h_{8}+\left(1-h_{8}\right)\Phi^{\beta^{\prime}}\right]{\rm Tr}\left(V_{\mu\nu}\right){\rm Tr}\left(V^{\mu\nu}\right)\ ,$
where $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}-i\left[V_{\mu},V_{\nu}\right]$, $V_{\mu}=\frac{1}{2}\left(g_{\omega}\omega_{\mu}+g_{\rho}\rho_{\mu}^{a}\tau^{a}\right)$. $m_{\rho}$, $m_{\omega}$ and $\mathcal{M}=m_{\pi}^{2}I_{2\times 2}$ are the masses of $\rho$, $\omega$ and $\pi$ mesons, respectively. The dilaton field $\chi$ is introduced as a nonlinear representation $\chi=f_{\chi}\Phi=f_{\chi}\exp\left(\sigma/f_{\chi}\right)$, and $\beta^{\prime}$ accounts for the anomalous dimension of gluon field operators, representing the deviation from the IR fixed point. According to Ref. Li and Ma (2017), $h_{i}$s, except $h_{5}$ and $h_{6}$, are chosen to be $1$ for simplification and the anomalous dimension $\gamma_{m}$ is simply taken to be $1$ (denoted as LOSS). $h_{5}$ and $h_{6}$ are constrained by the saddle point equations,
$4h_{5}+\left(4+\beta^{\prime}\right)h_{6}+2m_{\pi}^{2}f_{\pi}^{2}=0\ ,$
$12h_{5}+\left(4+\beta^{\prime}\right)\left(3+\beta^{\prime}\right)h_{6}+2m_{\pi}^{2}f_{\pi}^{2}=-m_{\sigma}^{2}f_{\chi}^{2}\ .$ (3)
In Lagrangian (II), pions are introduced as a nonlinear field $\xi=\sqrt{U}=e^{i\frac{\pi}{f_{\pi}}}$, where $\pi=\pi^{a}\tau^{a}$ with $a=1,2,3$ representing the isospin indices. Its covariant derivative is defined as
$D_{\mu}\xi=\left(\partial_{\mu}-iV_{\mu}\right)\xi=\left[\partial_{\mu}-i\frac{1}{2}\left(g_{\omega}\omega_{\mu}+g_{\rho}\rho_{\mu}^{a}\tau^{a}\right)\right]\xi\ .$ (4)
The Murrer-Cartan 1-forms $\hat{\alpha}_{\perp}^{\mu}$ and $\hat{\alpha}_{\|}^{\mu}$ are defined as
$\hat{\alpha}_{\perp,\|}^{\mu}=\frac{1}{2i}\left(D^{\mu}\xi\cdot\xi^{\dagger}\mp D^{\mu}\xi^{\dagger}\cdot\xi\right)\ .$ (5)
The baryonic part Lagrangian $\mathcal{L}_{B}$ is written as Li et al. (2017)
$\mathcal{L}_{B}$ $=$ $\left[g_{1}+(1-g_{1})\Phi^{\beta^{\prime}}\right]\bar{N}i\gamma_{\mu}D^{\mu}N-\left[g_{2}+\left(1-g_{2}\right)\Phi^{\beta^{\prime}}\right]m_{N}\Phi\bar{N}N$ (6)
${}+\left[g_{A}C_{A}+g_{A}\left(1-C_{A}\right)\Phi^{\beta^{\prime}}\right]\bar{N}\hat{\alpha}_{\perp}^{\mu}\gamma_{\mu}\gamma_{5}N+\left[g_{V_{\rho}}C_{V_{\rho}}+g_{V_{\rho}}\left(1-C_{V_{\rho}}\right)\Phi^{\beta^{\prime}}\right]\bar{N}\hat{\alpha}_{\|}^{\mu}\gamma_{\mu}N$
${}+\frac{1}{2}\left[g_{V_{0}}C_{V_{0}}+g_{V_{0}}\left(1-C_{V_{0}}\right)\Phi^{\beta^{\prime}}\right]\operatorname{Tr}\left[\hat{\alpha}_{\|}^{\mu}\right]\bar{N}\gamma_{\mu}N\ ,$
where $N$ is the iso-doublet of baryon field, and $g_{1}$ and $g_{2}$ are set to $1$ as suggested in Ref. Li and Ma (2017). For convenience, we introduce the combinations of the parameters:
$g_{\omega NN}=\frac{1}{2}\left(g_{V_{\rho}}+g_{V_{0}}-1\right)g_{\omega}\ ,$
$g_{\rho NN}=\frac{1}{2}\left(g_{V_{\rho}}-1\right)g_{\rho}\ ,$
$g^{SSB}_{\omega NN}=\frac{1}{2}\left[g_{V_{\rho}}\left(1-C_{V_{\rho}}\right)+g_{V_{0}}\left(1-C_{V_{0}}\right)\right]g_{\omega}\ ,$
$g^{SSB}_{\rho NN}=\frac{1}{2}\left[g_{V_{\rho}}\left(1-C_{V_{\rho}}\right)\right]g_{\rho}\ .$ (7)
By regarding the NM as homogeneous matter, the RMF approximation can be applied. Using Lagrangian (1), the EOMs of $\omega$ and $\rho$ can be obtained as
$m^{2}_{\omega}\Phi^{2}\omega-\left[g_{\omega NN}+g^{SSB}_{\omega NN}\left(\Phi^{\beta^{\prime}}-1\right)\right]\left(\rho_{n}+\rho_{p}\right)=0\ ,$
$m^{2}_{\rho}\Phi^{2}\rho-\left[g_{\rho NN}+g^{SSB}_{\rho NN}\left(\Phi^{\beta^{\prime}}-1\right)\right]\left(\rho_{p}-\rho_{n}\right)=0\ ,$ (8)
where $\rho$ and $\omega$ fields are denoted, respectively, as $\rho$ and $\omega$ for brevity. Similarly, the EOM of $\sigma$ field is derived as
$m^{2}_{\omega}\omega^{2}\Phi+m^{2}_{\rho}\rho^{2}\Phi$ $=$ $\frac{m^{4}_{N}\Phi^{3}}{\pi^{2}}\left[F\left(\frac{k_{p}}{m_{N}\Phi}\right)+F\left(\frac{k_{n}}{m_{N}\Phi}\right)\right]$ (9)
${}-2f^{2}_{\pi}m^{2}_{\pi}\Phi-4h_{5}\Phi^{3}-\left(4+\beta^{\prime}\right)h_{6}\Phi^{3+\beta^{\prime}}$
${}+g^{SSB}_{\omega NN}\beta^{\prime}\Phi^{\beta^{\prime}-1}\omega\left(\rho_{p}+\rho_{n}\right)+g^{SSB}_{\rho NN}\beta^{\prime}\Phi^{\beta^{\prime}-1}\rho\left(\rho_{p}-\rho_{n}\right)\ ,$
where $\frac{m_{N}^{3}\Phi^{3}}{\pi^{2}}F\left(\frac{k_{p(n)}}{m_{N}\Phi}\right)$ refers to scalar density $\langle\bar{p}p\rangle$ or $\langle\bar{n}n\rangle$ and $k_{p(n)}$ is the Fermi momentum of nucleons at zero temperature.
The energy density can be obtained via Lagrangian (1) with the solutions of EOMs mentioned above
$\langle\mathcal{H}\rangle$ $=$ ${-i\left\langle\bar{N}\gamma_{i}\partial^{i}N\right\rangle+m_{N}\Phi\langle\bar{N}N\rangle}+\frac{1}{2}m_{\omega}^{2}\Phi^{2}\omega^{2}+\frac{1}{2}m_{\rho}^{2}\Phi^{2}\rho^{2}-f_{\pi}^{2}m_{\pi}^{2}\Phi^{2}-h_{5}\Phi^{4}-h_{6}\Phi^{4+\beta^{\prime}}$ (10)
$=$ $\frac{m_{N}^{4}\Phi^{4}}{\pi^{2}}\left[f\left(\frac{k_{p}}{m_{N}\Phi}\right)+f\left(\frac{k_{n}}{m_{N}\Phi}\right)\right]+\frac{1}{2}m_{\omega}^{2}\Phi^{2}\omega^{2}+\frac{1}{2}m_{\rho}^{2}\Phi^{2}\rho^{2}$
${}-f_{\pi}^{2}m_{\pi}^{2}\Phi^{2}-h_{5}\Phi^{4}-h_{6}\Phi^{4+\beta^{\prime}}\ ,$
where $f\left(\frac{k_{p}(n)}{m_{N}\Phi}\right)=\int_{0}^{\frac{k_{p}(n)}{m_{N}}}x^{\prime 2}\sqrt{1+x^{\prime 2}}{\mathrm{d}}x^{\prime}$. It should be noted that the nonlinear Lagrangian (1) will lead to nonzero constant vacuum energy $\mathcal{E}_{0}=-f_{\pi}^{2}m_{\pi}^{2}-h_{5}-h_{6}$, which is neglected in our equation of state (EOS) calculations.
The above bsHLS (1) is constructed in a matter-free space. When applying it to a medium, it’s natural to expect that the parameters in the Lagrangian should be changed by medium, here density. We implement this density effect via Brown-Rho scaling (B-R scaling) Brown and Rho (1991, 2002), and refer to this density effect as intrinsic density dependence (IDD). Explicitly, the parameters in Lagrangian (1) scale as
$~{}\frac{m_{\rho(\omega,N)}^{*}}{m_{\rho(\omega,N)}}\approx\frac{f_{\pi}^{*}}{f_{\pi}}\approx\Phi^{*}\ ,\quad\frac{m_{\sigma}^{*}}{m_{\sigma}}\approx\left(\Phi^{*}\right)^{1+\frac{\beta^{\prime}}{2}}\ .$ (11)
A possible choice of $\Phi^{*}$ is $1/(1+r\frac{\rho_{n}+\rho_{p}}{n_{0}})$. Pion-nuclei bound state data Kienle and Yamazaki (2004) indicates $r\approx 0.2$, but we set it as a free parameter to fit the NM properties in this work. Chiral dynamics indicates that pion mass is not changed by medium, so we set $m_{\pi}^{*}/m_{\pi}\approx 1$. The saddle point equation (II) leads to
$h_{5}^{*}=$ 2($\frac{-2\left(2+\beta^{\prime}\right)m^{*2}_{\pi}f^{*2}_{\pi}+m^{*2}_{\sigma}f^{*2}_{\chi}}{4\beta^{\prime}}\ ,\quad h_{6}^{*}=$ $\frac{4m^{*2}_{\pi}f^{*2}_{\pi}-m^{*2}_{\sigma}f^{*2}_{\chi}}{\left(4+\beta^{\prime}\right)\beta^{\prime}}\ .$ (12)
After the above discussions, we finally obtain the energy density for phenomenological analysis as
$\mathcal{E}$ $=$ $\frac{m_{N}^{*4}\Phi^{4}}{\pi^{2}}\left[f\left(\frac{k_{p}}{m_{N}^{*}\Phi}\right)+f\left(\frac{k_{p}}{m_{N}^{*}\Phi}\right)\right]+\frac{1}{2}m_{\omega}^{*2}\Phi^{2}\omega^{2}+\frac{1}{2}m_{\rho}^{*2}\Phi^{2}\rho^{2}$ (13)
${}-f_{\pi}^{*2}m_{\pi}^{2}\Phi^{2}-h_{5}^{*}\Phi^{4}-h_{6}^{*}\Phi^{4+\beta^{\prime}}-\mathcal{E}_{0}^{*}\ .$
III Phenomenological analysis
In the phenomenological analysis, we choose vacuum values $f_{\pi}=92.4~{}\rm MeV$, $m_{N}=939~{}\rm MeV$, $m_{\omega}=783~{}\rm MeV$ and $m_{\rho}=765~{}\rm MeV$ Workman et al. (2022). The free parameters are $M_{\sigma}=m_{\sigma}f_{\chi}$, $\beta^{\prime}$, $r$, $g_{\omega NN}$, $g_{\rho NN}$, $g^{SSB}_{\omega NN}$ and $g^{SSB}_{\rho NN}$, which can be estimated by the properties of NM around saturation density $n_{0}$. The calculated results of NM properties are listed in Table 1, and the corresponding low energy constants (LECs) are given in Table 2. It can be seen that NM properties obtained from both bsHLS-L and bsHLS-H are consistent with empirical values.
Table 1: The properties of nuclear matter: $e_{0}$ is the binding energy of nucleon at $n_{0}$, $E_{\mathrm{sym}}(n)=\left.\frac{1}{2}\frac{\partial^{2}E(n,\alpha)}{\partial\alpha^{2}}\right|_{\alpha=0}$ is the symmetry energy, $\left.K_{0}=9n^{2}\frac{\partial^{2}E(n,0)}{\partial n^{2}}\right|_{n=n_{0}}$ is the incompressibility coefficient, $\left.J_{0}=27n^{3}\frac{\partial^{3}E(n,0)}{\partial n^{3}}\right|_{n=n_{0}}$ is the skewness coefficient and $L(n)=3n\frac{\partial E_{\mathrm{sym}}(n)}{\partial n}$ is the symmetry energy density slope. $n_{c}\approx 0.11\mathrm{fm^{-3}}$ is subsaturation cross density. Two sets of predictions are shown: bsHLS-L refers to the case, where the surge of $K(n)$ located at lower density regions, and bsHLS-H refers to the higher density case. $n_{0}$ is in the unit of $\rm fm^{-3}$, and the others are in the unit of $\rm MeV$.
Empirical bsHLS-L bsHLS-H
$n_{0}$ $0.155\pm 0.050$ Sedrakian et al. (2023) 0.159 0.159
$e_{0}$ $-15.0\pm 1.0$ Sedrakian et al. (2023) $-16.0$ $-16.0$
$K_{0}$ $230\pm 30$ Dutra et al. (2012) 232 284
$E_{\mathrm{sym}}(n_{c})$ $26.7\pm 0.2$ Zhang and Chen (2013) 20.8 20.9
$E_{\mathrm{sym}}(n_{0})$ $30.9\pm 1.9$ Lattimer and Lim (2013) 30.5 29.2
$E_{\mathrm{sym}}(2n_{0})$ $46.9\pm 10.1$ Li et al. (2019) 51.5 50.2
$L(n_{c})$ $43.7\pm 7.8$ Zhang and Chen (2014) 53.2 54.2
$L(n_{0})$ $52.5\pm 17.5$ Lattimer and Lim (2013) 85.9 68.3
$J_{0}$ $-700\pm 500$ Farine et al. (1997) $-767$ $-599$
Table 2: The estimation of parameters for bsHLS-L and bsHLS-H.
$M_{\sigma}(10^{5}\rm MeV^{2})$ $\beta^{\prime}$ $r$ $g_{\omega NN}$ $g_{\rho NN}$ $g^{SSB}_{\omega NN}$ $g^{SSB}_{\rho NN}$
bsHLS-L 1.05 0.395 0.161 11.5 3.78 16.3 9.45
bsHLS-H 2.30 1.15 0.191 11.0 4.17 8.85 4.85
To illustrate the features of bsHLS, we consider a Walecka-type model for comparison
$\mathcal{L}_{\rm RMF}$ $=$ $\bar{\psi}\left[i\gamma_{\mu}\partial^{\mu}-m_{N}-g_{\sigma}\sigma-g_{\omega_{NN}}\gamma_{\mu}\omega^{\mu}-g_{\rho_{NN}}\gamma_{\mu}\rho^{\mu a}\tau^{a}-g_{\delta}\delta^{a}\tau^{a}\right]\psi$ (14)
${}+\frac{1}{2}\left(\partial_{\mu}\sigma\partial^{\mu}\sigma-m_{\sigma}^{2}\sigma^{2}\right)-\frac{1}{3}g_{2}\sigma^{3}-\frac{1}{4}g_{3}\sigma^{4}$
${}-\frac{1}{2g^{2}}\operatorname{Tr}\left(V_{\mu\nu}V^{\mu\nu}\right)+\frac{1}{2}m_{\omega}^{2}\omega_{\mu}\omega^{\mu}+\frac{1}{4}c_{3}\left(\omega_{\mu}\omega^{\mu}\right)^{2}+\frac{1}{2}m_{\rho}^{2}\rho^{\mu a}\rho_{\mu}^{a}$
${}+\frac{1}{2}\Lambda_{V}\rho^{a}_{\mu}\rho^{\mu a}\omega_{\nu}\omega^{\nu}+\frac{1}{2}\left(\partial_{\mu}\delta^{a}\partial^{\mu}\delta^{a}-m^{2}_{\delta}\delta^{a}\delta^{a}\right)+\frac{1}{2}C_{\delta\sigma}\sigma^{2}\left(\delta^{a}\right)^{2}\ ,$
where $\sigma$, $\omega$, $\rho$ and $\delta$ mesons are introduced. The choices of parameters from some references are listed in Table 3.
Table 3: The choice of parameters for Walekca-type models in Eq. (14).
L-HS Horowitz and Serot (1981) NL1 Reinhard et al. (1986) TM1 Sugahara and Toki (1994) FSU-$\delta 6.7$ Li et al. (2022)
$n_{0}({\rm fm^{-3}})$ 0.149 0.152 0.145 0.148
$m_{N}$(MeV) 939 938 938 938
$m_{\sigma}$(MeV) 520 492 511 492
$m_{\omega}$(MeV) 783 795 783 783
$m_{\rho}$(MeV) 770 763 770 763
$g_{\sigma}$ 10.5 10.1 10.0 10.2
$g_{\omega NN}$ 13.8 13.3 12.6 13.4
$g_{\rho NN}$ 4.04 4.98 4.63 7.27
$g_{2}(\rm fm^{-1})$ 0 -12.2 -7.23 -8.09
$g_{3}$ 0 -36.3 0.618 5.88
$c_{3}$ 0 0 71.3 172
$g_{\delta}$ 0 0 0 6.70
$\Lambda_{V}$ 0 0 0 204
$m_{\delta}$ 0 0 0 980
$C_{\delta\sigma}$ 0 0 0 180
III.1 Nuclear matter properties
We first study the NM properties. From Table 1, one can see that, with appropriate choices of parameters in Table 2, our bsHLS can yield NM properties around (sub)saturation density, satisfying the constraints from empirical data.
(a) Symmetry energy as a function of density.
(b) Incompressibility as a function of density.
Figure 1: Incompressibility, $K=9\frac{{\rm d}P}{{\rm d}n}$, which is reduced to $K_{0}$ at $n_{0}$, and symmetry energy of symmetric NM from bsHLS and Walecka-type models.
Regarding the $L(n_{c})$ results listed in Table 1, both bsHLS-L and bsHLS-H exhibit the stiffness to align with the neutron skin thickness of $\rm Pb^{208}$ Reed et al. (2021), as well as the Walecka-type models compared in this work, shown in Fig. 1. Moreover, the $E_{\rm sym}$ behavior of bsHLS is quite similar to that of the compared Walecka-type models at low densities $n\lesssim n_{0}$. However, at intermediate densities, around $2n_{0}$, the models can be categorized into two sets: bsHLS and FSU-$\delta 6.7$ yield soft $E_{\rm sym}$, whereas L-HS, NL1 and TM1 give stiff one. The Walecka-type models without the $\delta$ meson are too stiff across the entire density range, while bsHLS provides a more reasonable behavior without introducing $\delta$. It will be seen later that this difference can affect the tidal deformation of NS.
For the incompressibility, the results of bsHLS and the compared Walekca-type models show significant differences at intermediate densities: Walecka-type models exhibit a simple behavior with density, whereas bsHLS shows a kink behavior around $(1\sim 2)n_{0}$. This kink behavior results in a peak structure in the sound velocity, and is attributed to manifestation of scale symmetry in nuclear matter Zhang et al. (2024).
III.2 Neutron star structures
Next, we study the NS structures using the EOSs discussed above. The results of NS structures are shown in Fig. 2 and Table 4.
Figure 2: The M-R relations from bsHLS and Walecka-type models. The constraints are estimated from Refs. Abbott et al. (2018, 2019); Riley et al. (2019, 2021)
Table 4: Tidal deformations from bsHLS and Walecka-type models defined in Eq. (14) and Table 3 for NS with mass of $1.4\ M_{\odot}$.
bsHLS-L bsHLS-H TM1 L-HS NL1 FSU-$\delta 6.7$
$\Lambda_{1.4}$ 2120 910 2240 2780 2620 878
From Fig. 2, one can see that the mass-radius (M-R) relation obtained from bsHLS-H falls within the constraints of astro-observations. Although the result from FSU-$\delta 6.7$ satisfies these constraints, it requires introducing a new degree of freedom $\delta$. The Walecka-type models without $\delta$ considered in this work cannot provide results within the constraints of NSs, though they reproduce the NM properties around $n_{0}$ Horowitz and Serot (1981); Reinhard et al. (1986); Sugahara and Toki (1994). This highlights the advantage of bsHLS in interpreting the structures of nuclei and NSs within a unified framework. It should be emphasized that the present analysis indicates that, in order to have realistic parameter space of nuclear model, including Walecka-type models, the data of both NM properties and NS structures should be considered in the statistical analysis simultaneously, as discussed in Ref. Guo et al. (2024). When meeting the NS constraints, the $M_{\rm max}$ of NSs predicted by these Walecka-type models is around $2M_{\odot}$, as shown in Ref. Özel and Freire (2016); Guo et al. (2024). This makes bsHLS more appealing, as the predicted $M_{\rm max}$ of bsHLS-H is $\sim 2.8M_{\odot}$. This predicted $M_{\rm max}$ may have profound implications for addressing the gap problem in the continuous mass distribution of supernova remnants Bailyn et al. (1998); Farr et al. (2011) and for understanding the gravitational wave event GW190814 Abbott et al. (2020).
For tidal deformations from Table 4, one can conclude that only the results of bsHLS-H and FSU-$\delta 6.7$ are close to the constraints of GW170817 Abbott et al. (2018). This is due to the softness of $E_{\rm sym}$ from these two models at intermediate densities, as shown in Fig. 1. Since $E_{\rm sym}$ from bsHLS-L is stiffer than bsHLS-H and FSU-$\delta 6.7$ below $\simeq 3n_{0}$, bsHLS-L gives larger $\Lambda_{1.4}$. The same reasoning applies to the Walecka-type models, NL1, L-HS, and TM1.
In summary, the above discussion indicates that bsHLS-H is a reasonable model for NM and NS. This implies that the scaling parameter $r\approx 0.19$, $\beta^{\prime}\approx 1.15$, $M_{\sigma}=m_{\sigma}f_{\chi}\approx 2.3\times 10^{5}~{}\rm MeV^{2}$ and the couplings between vector mesons and nucleons $g_{\rho NN}\approx 4.17$, $g_{\omega NN}\approx 11.0$. The scaling parameter $r\approx 0.19$ is consistent with pion-nuclei bound state data Kienle and Yamazaki (2004). The $\beta^{\prime}\approx 1.15$ agrees with estimates from skyrmion crystal approach Ma and Rho (2018); Shao and Ma (2022). If $f_{\chi}$ is taken to be $3f_{\pi}\approx 270~{}\rm MeV$, the power-counting mechanism of Crewther and Tuntall remains valid Crewther and Tunstall (2015), while $m_{\sigma}\approx 850~{}\rm MeV$. The coupling constant $g_{\rho NN}\approx 4.17$ aligns with the results from the OBE potential analysis of nucleon interactions Erkelenz (1974), and $g_{\omega NN}\approx 11.0$ is in agreement with the analyses of nucleon-nucleon scatterings Backman et al. (1985).
III.3 The patterns of scale symmetry and phenomenologies
The difference between bsHLS-L and bsHLS-H becomes significant in the M-R relation, as shown in Fig. 2, though it is not distinguishable from the NM properties around saturation density. The key difference lies in $\beta^{\prime}$, which connects to the existence of the pseudo-conformal structure at high densities Shao and Ma (2022). The $\beta^{\prime}$ of bsHLS-H aligns with the constraints of the pseudo-conformal limit, whereas the $\beta^{\prime}$ of bsHLS-L does not. This indicates the influence of scale symmetry on NS structures.
Moreover, since $\sigma$ is nonlinearly coupled with other mesons (see Eq. (II)) through a conformal compensator, its density dependence shows a kink behavior that is not observed in Walecka-type models Zhang et al. (2024). As the result, the order parameter $\langle\chi\rangle^{*}$, calculated based on bsHLS does not approach zero throughout the density regions, which is necessary for the (pseudo-)conformal limits of QCD at high densities Borsanyi et al. (2014); Ma and Rho (2019b); Shao and Ma (2022). To recover the expected behavior of $\langle\chi\rangle^{*}$, a possible approach is to couple an additional factor to $g_{\omega NN}$ to obtain the effective coupling $\tilde{g}_{\omega NN}=g_{\omega NN}/(1+R\frac{\rho_{p}+\rho_{n}}{n_{0}})$ Paeng et al. (2013); Zhang et al. (2024). It is found that, as listed in Table 5, the NM properties can be reproduced with the parameter set bsHLS-HS: $M_{\sigma}=2.25\times 10^{5}\ \rm MeV^{2}$, $\beta^{\prime}=1.14$, $g_{\omega NN}=11.5$, $g_{\rho NN}=4.27$, $g^{SSB}_{\omega NN}=8.70$, $g^{SSB}_{\rho NN}=4.85$, $r=0.20$ and $R=0.02$.
Table 5: The quantities of nuclear matter with additional suppressions of $\tilde{g}_{\omega NN}$, and the definitions and constraints are the same as Table 1.
$n_{0}$ $e_{0}$ $K_{0}$ $E_{\mathrm{sym}}(n_{c})$ $E_{\mathrm{sym}}(n_{0})$ $E_{\mathrm{sym}}(2n_{0})$ $L(n_{c})$ $L(n_{0})$ $J_{0}$
bsHLS-HS 0.159 -16.0 259 21.6 30.3 52.9 55.4 74.5 -720
And, as shown in Fig. 3, the M-R relations are roughly consistent with observational constraints, but there is a decrease of $M_{\rm max}$ compared to bsHLS-H.
Figure 3: NS structure results with/without $g_{\omega NN}$ suppression.
In order to understand the discrepancies in the M-R relations, the incompressibilities and symmetry energies are also calculated for bsHLS-H and bsHLS-HS, shown in Fig. 4.
(a) Imcompressibilities as a function of density in symmetric NM.
(b) Symmetry energies as a function of density.
Figure 4: NM property results with/without $G_{\omega}$ suppression.
It can be seen that the symmetry energy remains almost the same in both cases. However, the incompressibility of bsHLS-HS is much softer than that of bsHLS-H at high densities, leading to a decrease in $M_{\rm max}$. At low or intermediate densities, the incompressibilities of both cases are similar, which results in the M-R relations being comparable in the region of NS constraints.
Additionally, the NM properties were investigated with B-R scaling turned off. We find that with the parameter set bsHLS-N:$M_{\sigma}=1.11\times 10^{5}\ \rm MeV^{2}$, $\beta^{\prime}=1.10$, $g_{\omega NN}=-11.0$, $g_{\rho NN}=4.17$, $g^{SSB}_{\omega NN}=-192$ and $g^{SSB}_{\rho NN}=4.85$, the NM properties can still be reproduced (e.g., $n_{0}=0.16\ \rm{fm}^{-3}$, $e_{0}=-16.0\ \rm{MeV}$, $K_{0}=241\ \rm{MeV}$, $E_{{\rm sym}}(n_{0})=29.3\ \rm{MeV}$). However, this approach results in a much softer scale symmetry parameter $\langle\chi\rangle^{*}$ flow compared to the cases with B-R scaling. As a result, the M-R relation becomes unnatural and deviates from the constraints, as shown in Fig. 5.
(a) $\langle\chi\rangle^{*}/f_{\chi}$ as a function of density.
(b) MR relation of NS.
Figure 5: The $\langle\chi\rangle^{*}/f_{\chi}$ in pure neutron matter and NS structure results with/without B-R scaling.
All these findings, including the necessity of introducing $\tilde{g}_{\omega NN}$ suppression and B-R scaling, signify the importance of properly parametrizing the scale symmetry in densities. This reminds us of the past work on the quenching factors of $g_{A}$ which affect the $\beta$ decay of neutrons in dense environments (see Ref. Ma and Rho (2020b) for details). From the comparison among the various cases discussed above, valuable lessons about NS and NM properties have been learned, providing guidance on hadron interaction parameterization to describe dense NM. Since the differences between bsHLS and Walecka-type models in our analysis are at intermediate density regions, not far from $n_{0}$, it is expected that examining bsHLS in dense systems will be promising in future experiments.
IV Conclusion and discussion
In this work, the bsHLS, constructed based on the philosophy of $\chi\rm EFT$ with HLS and a possible IR fixed point of QCD at low energies, is applied to dense environments using the RMF approximation, with free parameters fixed by pinning the nuclei structure data around $n_{0}$. The NM properties and NS structures can be well reproduced, closely matching empirical values in bsHLS-H case.
Without introducing many freedoms, such as the $\delta$ meson, and with operators organized to respect chiral and scale symmetry considerations and expanded by chiral-scale orders, the bsHLS can provide a reasonable behavior of NM properties from subsaturation to intermediate densities, e.g., $K(n)$ and $E_{\rm sym}(n)$, compared to Walecka-type models. The NS structures are sensitive to NM properties in these density regions, making bsHLS outperform Walecka-type models in describing a wider range of densities. More specifically, the kink behavior of the $\sigma$ field in bsHLS at intermediate densities allows the $M_{\rm max}$ to reach nearly $3\ M_{\odot}$, while other NS observational constraints are still statisfied.
Besides, the behaviors of symmetry patterns in dense environments are also found to be pivotal to macroscopic phenomena: If there is no restoration point of scale symmetry at certain densities, such as the $\beta^{\prime}$ value of bsHLS-L, the NS structures will fall outside observational constraints; The $M_{\rm max}$ of predicted NSs is influenced by the behavior of order parameter of scale symmetry, $\langle\chi\rangle^{*}$. Furthermore, the study on the flow of $\langle\chi\rangle^{*}$ with densities suggests the necessity of introducing an additional suppression factor for $\tilde{g}_{\omega NN}$ which could be an interesting problem for further investigation.
In summary, introducing bsHLS to NM studies is a promising approach due to its close relation to QCD symmetry patterns and the effective potentials organized by chiral-scale orders, which have already proven successful in describing scattering experiments at vacuum. Furthermore, the difference between bsHLS and Walecka-type models is not far from $n_{0}$, making it possible to be verified in future experiments, such as heavy ion collisions. The relationship between microscopic symmetries and macroscopic phenomena found in this work is also a valuable topic for further study.
Acknowledgments
The work of Y. L. M. is supported in part by the National Science Foundation of China (NSFC) under Grant No. 12347103 and No. 11875147.

### v3

@@
 Utilizing these models, NM properties can be calculated using the relativistic mean field (RMF) approach Walecka (1974); Wiringa et al.
-(1988), which is the most practical and economical framework to introduce density effects.
-It is recognized that Walecka-type models lack considerations of QCD symmetry patterns, the valid region of effective operators, and theoretical errors.
+(1988); Gao et al.
+(2022); Gao and Harada (2024), which is the most practical and economical framework to introduce density effects.
+It is recognized that Walecka-type models lack the consideration of QCD symmetry patterns, the valid region of effective operators, and theoretical errors.
 Chiral effective field theory ($\rm{\chi EFT}$), thanks to the pioneering works by Weinberg Weinberg (1979, 1991), offers a powerful framework for studying nuclear forces at long ranges, anchored on QCD symmetry.
@@
 (2007), taking the “leading order scale symmetry (LOSS)” approximation where the trace anomaly effect enters only through the dilaton potential, which breaks scale symmetry explicitly and spontaneously, we found that the chiral-scale EFT with few parameters can successfully describe not only NM at the saturation density but also the compact-star matter at $n=(5-7)n_{0}$ Ma and Rho (2020a).
-A novel phenomena, which was not realized, is that in nuclear matter at densities $n=(2-4)n_{0}$ the sound velocity of NM saturates the conformal limit $v_{s}^{2}=1/3$, but the trace of the energy-momentum tensor does not vanish, that is, the NM exhibits a pseudo-conformal structure Ma and Rho (2019a, b); Ma et al.
+A novel phenomena, which was not realized before, is that in nuclear matter at densities $n=(2-4)n_{0}$ the sound velocity of NM saturates the conformal limit $v_{s}^{2}=1/3$, but the trace of the energy-momentum tensor does not vanish, that is, the NM exhibits a pseudo-conformal structure Ma and Rho (2019a, b); Ma et al.
 (2019).
+And the corrections to LOSS in the baryonic part have been found crucial for understanding the quenched $g_{A}$ value in the super-allowed Gamow-Teller transitions of heavy nuclei Lubos et al.
+(2019); Ma and Rho (2020b).
 In this work, we study NM properties using bsHLS with the RMF method.
-The corrections to LOSS in the baryonic part, which have been found crucial for understanding the quenched $g_{A}$ value in the super-allowed Gamow-Teller transitions of heavy nuclei Lubos et al.
-(2019); Ma and Rho (2020b).
 After identifying the NM properties around $n_{0}$, it was found that bsHLS using the RMF method can yield reasonable results.
@@
 (2018, 2019) and PSR J0740+6620 Miller et al.
+(2021); Riley et al.
 (2021), whereas Walecka-type models can only reach $2M_{\odot}$ within these constraints, as analyzed in Ref.
@@
 II bsHLS in nuclear medium
-In this section, we establish the theoretical framework of bsHLS and the EoS of NM from the RMF approach.
-Here, we only interest in NM made of nucleons, we reduce the bsHLS Lagrangian to the two-flavor case, $u$ and $d$.
+In this section, we establish the theoretical framework of bsHLS and derive the equation of state (EoS) for NM within the relativistic mean field (RMF) approach.
+Focusing on NM only composed of nucleons, we restrict the bsHLS Lagrangian to the two-flavor case, involving only $u$ and $d$ quarks.
 The leading order bsHLS Lagrangian $\mathcal{L}$ can be decomposed into the baryonic part $\mathcal{L}_{B}$ and the mesonic part $\mathcal{L}_{M}$:
@@
 where $N$ is the iso-doublet of baryon field, and $g_{1}$ and $g_{2}$ are set to $1$ as suggested in Ref. Li and Ma (2017). For convenience, we introduce the combinations of the parameters:
-$g_{\omega NN}=\frac{1}{2}\left(g_{V_{\rho}}+g_{V_{0}}-1\right)g_{\omega}\ ,$
-$g_{\rho NN}=\frac{1}{2}\left(g_{V_{\rho}}-1\right)g_{\rho}\ ,$
+$g_{\omega NN}=\frac{1}{2}\left(g_{V_{p}}C_{V_{p}}+g_{V_{0}}C_{V_{0}}-1\right)g_{\omega}\ ,$
+$g_{\rho NN}=\frac{1}{2}\left(g_{V_{p}}C_{V_{p}}-1\right)g_{\rho}\ ,$
 $g^{SSB}_{\omega NN}=\frac{1}{2}\left[g_{V_{\rho}}\left(1-C_{V_{\rho}}\right)+g_{V_{0}}\left(1-C_{V_{0}}\right)\right]g_{\omega}\ ,$
@@
 III Phenomenological analysis
-In the phenomenological analysis, we choose vacuum values $f_{\pi}=92.4~{}\rm MeV$, $m_{N}=939~{}\rm MeV$, $m_{\omega}=783~{}\rm MeV$ and $m_{\rho}=765~{}\rm MeV$ Workman et al.
+In the phenomenological analysis, we choose vacuum values $f_{\pi}=92.4~{}\rm MeV$, $m_{N}=939~{}\rm MeV$, $m_{\pi}=140~{}\rm MeV$, $m_{\omega}=783~{}\rm MeV$ and $m_{\rho}=765~{}\rm MeV$ Workman et al.
 (2022).
@@
 $K_{0}$ $230\pm 30$ Dutra et al. (2012) 232 284
-$E_{\mathrm{sym}}(n_{c})$ $26.7\pm 0.2$ Zhang and Chen (2013) 20.8 20.9
+$E_{\mathrm{sym}}(n_{c})$ $22.4\pm 2.3$ Chen (2011); Liu et al. (2010) 20.8 20.9
 $E_{\mathrm{sym}}(n_{0})$ $30.9\pm 1.9$ Lattimer and Lim (2013) 30.5 29.2
@@
 $\Lambda_{V}$ 0 0 0 204
-$m_{\delta}$ 0 0 0 980
+$m_{\delta}$(MeV) 0 0 0 980
 $C_{\delta\sigma}$ 0 0 0 180
 III.1 Nuclear matter properties
-We first study the NM properties.
+The properties of NM is analyzed at first.
 From Table 1, one can see that, with appropriate choices of parameters in Table 2, our bsHLS can yield NM properties around (sub)saturation density, satisfying the constraints from empirical data.
@@
 III.2 Neutron star structures
-Next, we study the NS structures using the EOSs discussed above. The results of NS structures are shown in Fig. 2 and Table 4.
-Figure 2: The M-R relations from bsHLS and Walecka-type models. The constraints are estimated from Refs. Abbott et al. (2018, 2019); Riley et al. (2019, 2021)
-Table 4: Tidal deformations from bsHLS and Walecka-type models defined in Eq. (14) and Table 3 for NS with mass of $1.4\ M_{\odot}$.
+Next, the NS structures are studied using the EOSs discussed above for pure neutron matter (PNM). The results of NS structures are shown in Fig. 2 and Table 4.
+Figure 2: The M-R relations from bsHLS and Walecka-type models.
+The constraints are estimated from Refs.
+Abbott et al.
+(2018, 2019); Riley et al.
+(2019, 2021).
+The M-R relation is calculated by solving the Tolman-Oppenheimer-Volkoff (TOV) equation Oppenheimer and Volkoff (1939); Tolman (1939).
+All EoSs for TOV equation calculation are interpolated to a BPS EoS Baym et al.
+(1971) below $0.01\ \rm fm^{-3}$ Arnett and Bowers (1977).
+Table 4: Tidal deformations from bsHLS and Walecka-type models defined in Eq.
+(14) and Table 3 for NS with mass of $1.4\ M_{\odot}$.
+They are calculated from the EOSs used for the M-R relation in Fig.
+2, using the formalisms in Ref.
+Postnikov et al.
+(2010).
 bsHLS-L bsHLS-H TM1 L-HS NL1 FSU-$\delta 6.7$
@@
 (2024).
-When meeting the NS constraints, the $M_{\rm max}$ of NSs predicted by these Walecka-type models is around $2M_{\odot}$, as shown in Ref.
+When meeting the NS constraints, the $M_{\rm max}$ of NSs predicted by these Walecka-type models and chiral nuclear force models is around or slightly above $2M_{\odot}$, as discussed in Refs.
 Özel and Freire (2016); Guo et al.
+(2024); Fan et al.
 (2024).
@@
 In summary, the above discussion indicates that bsHLS-H is a reasonable model for NM and NS.
-This implies that the scaling parameter $r\approx 0.19$, $\beta^{\prime}\approx 1.15$, $M_{\sigma}=m_{\sigma}f_{\chi}\approx 2.3\times 10^{5}~{}\rm MeV^{2}$ and the couplings between vector mesons and nucleons $g_{\rho NN}\approx 4.17$, $g_{\omega NN}\approx 11.0$.
+It implies that the scaling parameter $r\approx 0.19$, $\beta^{\prime}\approx 1.15$, $M_{\sigma}=m_{\sigma}f_{\chi}\approx 2.3\times 10^{5}~{}\rm MeV^{2}$, and the couplings between vector mesons and nucleons $g_{\rho NN}\approx 4.17$, $g_{\omega NN}\approx 11.0$.
 The scaling parameter $r\approx 0.19$ is consistent with pion-nuclei bound state data Kienle and Yamazaki (2004).
-The $\beta^{\prime}\approx 1.15$ agrees with estimates from skyrmion crystal approach Ma and Rho (2018); Shao and Ma (2022).
-If $f_{\chi}$ is taken to be $3f_{\pi}\approx 270~{}\rm MeV$, the power-counting mechanism of Crewther and Tuntall remains valid Crewther and Tunstall (2015), while $m_{\sigma}\approx 850~{}\rm MeV$.
+And the $\beta^{\prime}\approx 1.15$ agrees with estimates from skyrmion crystal approach Ma and Rho (2018); Shao and Ma (2022).
+If $f_{\chi}$ is taken to be $3f_{\pi}\approx 270~{}\rm MeV$, $m_{\sigma}\approx 850~{}\rm MeV$, while the power-counting mechanism of Crewther and Tuntall remains valid Crewther and Tunstall (2015).
 The coupling constant $g_{\rho NN}\approx 4.17$ aligns with the results from the OBE potential analysis of nucleon interactions Erkelenz (1974), and $g_{\omega NN}\approx 11.0$ is in agreement with the analyses of nucleon-nucleon scatterings Backman et al.
@@
 2, though it is not distinguishable from the NM properties around saturation density.
-The key difference lies in $\beta^{\prime}$, which connects to the existence of the pseudo-conformal structure at high densities Shao and Ma (2022).
+The key difference lies in $\beta^{\prime}$, which is connected to the existence of the pseudo-conformal structure at high densities Shao and Ma (2022).
 The $\beta^{\prime}$ of bsHLS-H aligns with the constraints of the pseudo-conformal limit, whereas the $\beta^{\prime}$ of bsHLS-L does not.
-This indicates the influence of scale symmetry on NS structures.
+This highlights the impact of scale symmetry patterns on NS structures.
 Moreover, since $\sigma$ is nonlinearly coupled with other mesons (see Eq.
@@
 bsHLS-HS 0.159 -16.0 259 21.6 30.3 52.9 55.4 74.5 -720
-And, as shown in Fig. 3, the M-R relations are roughly consistent with observational constraints, but there is a decrease of $M_{\rm max}$ compared to bsHLS-H.
+And, as shown in Fig. 3, the M-R relations for PNM are roughly consistent with observational constraints, but there is a decrease of $M_{\rm max}$ compared to bsHLS-H.
 Figure 3: NS structure results with/without $g_{\omega NN}$ suppression.
@@
 At low or intermediate densities, the incompressibilities of both cases are similar, which results in the M-R relations being comparable in the region of NS constraints.
-Additionally, the NM properties were investigated with B-R scaling turned off.
+Additionally, the NM properties are investigated with B-R scaling turned off.
 We find that with the parameter set bsHLS-N: $M_{\sigma}=1.11\times 10^{5}\ \rm MeV^{2}$, $\beta^{\prime}=1.10$, $g_{\omega NN}=-11.0$, $g_{\rho NN}=4.17$, $g^{SSB}_{\omega NN}=-192$ and $g^{SSB}_{\rho NN}=4.85$, the NM properties can still be reproduced (e.g., $n_{0}=0.16\ \rm{fm}^{-3}$, $e_{0}=-16.0\ \rm{MeV}$, $K_{0}=241\ \rm{MeV}$, $E_{{\rm sym}}(n_{0})=29.3\ \rm{MeV}$).
@@
 Without introducing many freedoms, such as the $\delta$ meson, and with operators organized to respect chiral and scale symmetry considerations and expanded by chiral-scale orders, the bsHLS can provide a reasonable behavior of NM properties from subsaturation to intermediate densities, e.g., $K(n)$ and $E_{\rm sym}(n)$, compared to Walecka-type models.
-The NS structures are sensitive to NM properties in these density regions, making bsHLS outperform Walecka-type models in describing a wider range of densities.
-More specifically, the kink behavior of the $\sigma$ field in bsHLS at intermediate densities allows the $M_{\rm max}$ to reach nearly $3\ M_{\odot}$, while other NS observational constraints are still statisfied.
+And the NS structures are sensitive to NM properties at these density regions, making bsHLS outperform Walecka-type models in describing a wider range of densities.
+More specifically, the kink behavior of the $\sigma$ field in bsHLS at intermediate densities allows the $M_{\rm max}$ to reach nearly $3\ M_{\odot}$ for PNM, while other NS observational constraints are still statisfied.
 Besides, the behaviors of symmetry patterns in dense environments are also found to be pivotal to macroscopic phenomena: If there is no restoration point of scale symmetry at certain densities, such as the $\beta^{\prime}$ value of bsHLS-L, the NS structures will fall outside observational constraints; The $M_{\rm max}$ of predicted NSs is influenced by the behavior of order parameter of scale symmetry, $\langle\chi\rangle^{*}$.
-Furthermore, the study on the flow of $\langle\chi\rangle^{*}$ with densities suggests the necessity of introducing an additional suppression factor for $\tilde{g}_{\omega NN}$ which could be an interesting problem for further investigation.
+Furthermore, the study on the flow of $\langle\chi\rangle^{*}$ with densities suggests the necessity of introducing an additional suppression factor for $\tilde{g}_{\omega NN}$, and it could be an interesting problem for further investigation.
 In summary, introducing bsHLS to NM studies is a promising approach due to its close relation to QCD symmetry patterns and the effective potentials organized by chiral-scale orders, which have already proven successful in describing scattering experiments at vacuum.
 Furthermore, the difference between bsHLS and Walecka-type models is not far from $n_{0}$, making it possible to be verified in future experiments, such as heavy ion collisions.
-The relationship between microscopic symmetries and macroscopic phenomena found in this work is also a valuable topic for further study.
+The relationship between microscopic symmetries and macroscopic phenomena found in this work is also a valuable topic to be further studied.
 Acknowledgments
+We would like to thank Yi-Zhong Fan for his valuable comments.
 The work of Y. L. M. is supported in part by the National Science Foundation of China (NSFC) under Grant No. 12347103 and No. 11875147.